# HKAN: Hierarchical Kolmogorov-Arnold Network without Backpropagation

### v1

@@ Abstract
This paper introduces the Hierarchical Kolmogorov-Arnold Network (HKAN), a novel network architecture that offers a competitive alternative to the recently proposed Kolmogorov-Arnold Network (KAN). Unlike KAN, which relies on backpropagation, HKAN adopts a randomized learning approach, where the parameters of its basis functions are fixed, and linear aggregations are optimized using least-squares regression. HKAN utilizes a hierarchical multi-stacking framework, with each layer refining the predictions from the previous one by solving a series of linear regression problems. This non-iterative training method simplifies computation and eliminates sensitivity to local minima in the loss function. Empirical results show that HKAN delivers comparable, if not superior, accuracy and stability relative to KAN across various regression tasks, while also providing insights into variable importance. The proposed approach seamlessly integrates theoretical insights with practical applications, presenting a robust and efficient alternative for neural network modeling.
Index Terms:
Kolmogorov-Arnold network, neural networks, multi-stacking, randomized learning.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Kolmogorov-Arnold Networks (KANs), introduced in [1], represent a paradigm shift in neural network (NN) architecture, offering a promising alternative to traditional Multi-Layer Perceptrons (MLPs). Rooted in the Kolmogorov-Arnold representation theorem, which states that any multivariate continuous function $f:[0,1]^{n}\to\mathbb{R}$ can be represented as a superposition of continuous functions of a single variable and the binary operation of addition, KANs fundamentally reimagine the structure of NNs.
KANs introduce a significant architectural innovation: unlike MLPs with fixed activation functions on nodes, KANs employ learnable activation functions on edges. Notably, KANs eliminate linear weights entirely, replacing each weight parameter with a univariate function parametrized as a spline. This seemingly simple modification yields significant improvements in both accuracy and interpretability. In terms of accuracy, smaller KAN models consistently achieve comparable or superior performance to larger MLPs in data fitting and partial differential equations solving tasks. Both theoretical analysis and empirical evidence suggest that KANs exhibit faster neural scaling laws than MLPs. Regarding interpretability, KANs offer intuitive visualization and facilitate easy interaction with human users, enhancing their potential as collaborative tools for scientific discovery.
@@ I-A Related Work
Despite these advancements, KANs are not without criticism. Some studies argue that KAN outperforms MLPs primarily in symbolic formula representation but falls short in tasks like computer vision, natural language processing, and audio processing [8]. Tran et al. [9] reported that despite their theoretical advantages, KANs do not consistently outperform MLPs in practical classification tasks. Additionally, their hardware implementations tend to be less efficient, with higher resource usage and latency. Sensitivity to noise is another limitation; even minimal noise in the data can significantly degrade performance [10].
@@
The flexibility of KANs has led to explorations with various activation functions beyond B-splines, including wavelets [22], radial basis functions [23], Fourier series [24], Jacobi basis functions [25], rational functions [26], and ReLU [27]. A comprehensive comparison of activation functions used in KAN architectures is available in [28].
@@ I-B Motivation and Contributions
KAN models are traditionally trained using backpropagation algorithms, which rely on gradients of the network’s loss function with respect to its parameters. However, gradient-based learning processes are sensitive to issues such as local minima, flat regions, and saddle points in the loss function. Additionally, gradient calculations can be computationally expensive, particularly for deep and wide network architectures, complex target functions, and large training datasets.
@@
Our approach begins with utilizing fixed parameters for basis functions, determined either by data or randomly. These basis functions are then combined in multiple blocks using linear regression. The resulting block functions (activation functions) are subsequently combined through linear regression, and this iterative process is repeated across subsequent layers to form higher-level representations. Combining diverse blocks corresponds to ensembling, while performing it layer by layer constitutes a multi-stacking approach. This hierarchical modeling of the target function progressively enhances accuracy at each level, eliminating the need for backpropagation.
@@
    Multi-Stacking Approach for Prediction: Our hierarchical KAN (HKAN) implements hierarchically multi-stacking approach to built predictions. In each layer, meta-learners combine predictions performed by weak learners (univariate models). Subsequent layers, fed by predictions from previous layers, successively refine the results, enhancing overall accuracy layer by layer.
@@
    Empirical Results for Regression Problems: We provide comprehensive empirical evidence demonstrating that our HKAN outperforms standard KAN in a range of regression problems.
@@ II Preliminary
II-A Kolmogorov–Arnold Representation Theorem
The Kolmogorov–Arnold representation theorem, also known as the superposition theorem, stands as a cornerstone in the theory of function approximation. This profound result asserts that any continuous function of several variables can be represented as a composition of continuous functions of one variable and addition.
@@ II-A Kolmogorov–Arnold Representation Theorem
The Kolmogorov–Arnold representation theorem stands as a bridge between pure mathematics and applied computational science, highlighting the potential for representing complex functions through simpler components while also illustrating the challenges in translating theoretical results into practical applications.
II-B Kolmogorov–Arnold Networks (KANs)
Paper [1] extends and modifies the Kolmogorov-Arnold representation theorem to create Kolmogorov-Arnold Networks (KANs) in several key ways. While the original theorem uses a 2-layer network with a specific width in the hidden layer, KANs generalize this to allow arbitrary widths and depths, stacking multiple ”KAN layers”.
A KAN layer is defined as a matrix of activation functions $\phi_{q,p}$, where $q$ is not restricted to the theoretical limit of $12n+$ 1:
@@ II-B Kolmogorov–Arnold Networks (KANs)
where $x_{p}$ denotes the $p$-th input to the layer, $n_{in}$ denotes the number of inputs, and $n_{out}$ denotes the number of outputs (not restricted to $12n+$ 1).
@@
Furthermore, the paper provides tools for users to visualize and modify KANs, making them more interpretable and interactive. This approach takes the core idea of representing multivariate functions using univariate functions and addition, and extends it into a flexible, trainable NN architecture with theoretical guarantees and practical advantages over standard MLPs. In essence, [1] modernizes the Kolmogorov-Arnold theory for use in contemporary machine learning, offering a new perspective on function approximation and NN design.
@@ III Hierarchical KAN
$N$ number of training samples
@@
III-A Architecture
HKAN is an advanced NN architecture inspired by the Kolmogorov-Arnold representation theorem. While it shares similarities with the KAN architecture, HKAN introduces unique components and a distinct training process. The architecture of HKAN is shown in Fig. 1.
@@ III-A Architecture
Let $\mathbf{z}^{(l-1)}$ be the input vector for layer $l$. The first layer of HKAN takes input vector $\mathbf{z}^{(0)}=\mathbf{x}=[x_{1},...,x_{n}]^{\top}\in\mathbb{R}^{n}$. Each component of the input vector is transformed by a group of blocks. Within each group, there are $n^{(l)}$ blocks, and each block projects its input nonlinearly using $m^{(l)}$ basis functions (BaFs). Consider two common basis functions: Gaussian
@@
The configuration of BaFs plays a crucial role in the network’s performance. The number of BaFs, together with the smoothing parameter, serve as hyperparameters that define the block’s flexibility and balance the trade-off between variance and bias in the output. The locations of the BaFs, denoted as $\mu^{(l)}_{q,p,r}$ for the $r$-th function in the block corresponding to $\phi^{(l)}_{q,p}$, define the position of the maximum for Gaussian functions or the inflection point for sigmoid functions. To distribute BaFs in a block across the input interval (typically a bounded region of $[0,1]$), these locations can be selected in two ways:
@@
The support point method aligns the BaFs with the data distribution, avoiding empty regions in the input space (see Fig. 2).
Figure 2: Illustration of block function composition using the support point method. Red markers represent the support points that determine the placement of BaFs.
The BaFs within a block, $g^{(l)}_{q,p,r}:\mathbb{R}\rightarrow\mathbb{R}$, are combined using linear regression:
Resulting function $\phi^{(l)}_{q,p}:\mathbb{R}^{m^{(l)}}\rightarrow\mathbb{R}$ is called a block function (BlF). The weights of each BlF, $c^{(l)}_{q,p,r}$, are determined using least squares by minimizing the sum of squared residuals:
@@
where $\mathbf{c}^{(l)}_{q,p}=[c^{(l)}_{q,p,1},...,c^{(l)}_{q,p,m^{(l)}}]^{\top}$, $\mathbf{y}=[y_{1},...,y_{N}]^{\top}$ and $\mathbf{G}^{(l)+}_{q,p}$ is the Moore–Penrose generalized inverse of the BaF response matrix to the $N$ training data points or their projections:
@@
For the $p$-th input, each block in a group fits different function (due to random positioning of BaFs) that approximate the target function based on this input. Consequently, there are a total of $n^{(l-1)}n^{(l)}$ BlFs in layer $l$. In the subsequent step, these BlFs are combined linearly by $n^{(l)}$ $h$-functions. Each $q$-th $h$-function takes as input the $q$-th BlF from every group. Thus, each $h$-function approximates the target function based on different projections of all components of the input pattern for layer $l$:
The weights of this combination are calculated as:
@@
where $\mathbf{w}^{(l)}_{q}=[w^{(l)}_{q,1},...,w^{(l)}_{q,n^{(l)}}]^{\top}$, and $\mathbf{\Phi}^{(l)+}_{q}$ is the Moore–Penrose generalized inverse of the BlF response matrix to the projections of $N$ training data points:
@@
The output of layer $l$ is given by $\mathbf{z}^{(l)}=\hat{\mathbf{y}}^{(l)}=[\hat{y}^{(l)}_{1},...,\hat{y}^{(l)}_{n^{(l)}}]^{\top}\in\mathbb{R}^{n^{(l)}}$, where each component $\hat{y}^{(l)}_{q}=h^{(l)}_{q}(\mathbf{z}^{(l-1)})$. This output is then fed to the next layer and processed using (5)-(11).
Following the design of the original KAN, the structure of the top layer (layer $L$) in HKAN differs slightly from the preceding layers, as it includes only one BaF per input. The final output of HKAN is obtained through a linear combination of the $n^{(L)}=n^{(L-1)}$ BlFs:
$h^{(L)}(\mathbf{z}^{(L-1)})=\sum_{p=1}^{n^{(L)}}w^{(L)}_{p}\phi^{(L)}_{q}(z^{(L-1)}_{p})$ (12)

### v2

@@ Abstract
This paper introduces the Hierarchical Kolmogorov-Arnold Network (HKAN), a novel neural model that eliminates the need for backpropagation. While structurally related to the standard Kolmogorov-Arnold Network (KAN), HKAN employs a randomized learning framework and a hierarchical multi-stacking design, where each layer refines the approximations produced by the preceding one through a sequence of convex optimization subproblems. This non-iterative training strategy ensures high computational efficiency and numerical stability while preserving strong approximation accuracy. Experimental results on both synthetic and real-world regression tasks show that HKAN achieves accuracy comparable to or exceeding that of standard KANs and Multi-Layer Perceptrons, while reducing training time. Moreover, HKAN enhances interpretability by incorporating a built-in mechanism for assessing the importance of input variables. The proposed framework thus bridges theoretical rigor and practical utility, offering a robust, transparent, and computationally efficient alternative to gradient-based neural models.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
Kolmogorov-Arnold Networks (KANs) represent a paradigm shift in neural network (NN) architecture, offering a promising alternative to traditional Multi-Layer Perceptrons (MLPs). Rooted in the Kolmogorov-Arnold representation theorem (KART), which states that any multivariate continuous function $f:[0,1]^{n}\to\mathbb{R}$ can be represented as a superposition of continuous functions of a single variable and the binary operation of addition, KANs fundamentally redefine the structure of neural computation.
In this study, we refer specifically to the KAN model introduced by Liu et al. [1], which sparked the recent surge of research activity in this field. However, the conceptual origins of KANs date back to the 1980s. To the best of our knowledge, an architecture conceptually similar to that of Liu et al., incorporating spline-based basis functions, was first proposed by Igelnik and Parikh in 2003 [2].
KANs introduce a significant architectural innovation: unlike MLPs with fixed activation functions on nodes, KANs employ learnable activation functions on edges. Notably, KANs eliminate linear weights entirely, replacing each weight parameter with a univariate function – specifically, a spline, as in [1]. This seemingly simple modification yields significant improvements in both accuracy and interpretability. In terms of accuracy, smaller KAN models consistently achieve comparable or superior performance to larger MLPs in data fitting and partial differential equations solving tasks. Both theoretical analysis and empirical evidence suggest that KANs exhibit faster neural scaling laws than MLPs. Regarding interpretability, KANs offer intuitive visualization and facilitate easy interaction with human users, enhancing their potential as collaborative tools for scientific discovery.
@@ I-A Related Work
The origins of KANs can be traced back to the 1980s. Hecht-Nielsen was the first to recognize that KART could be applied in NN computing. He demonstrated that Kolmogorov’s superposition theorem could be interpreted as a four-layer feedforward NN [3]. Kurkova later observed that the fixed number of basis functions in KART could be replaced by a variable number, thereby transforming the task of function representation into one of function approximation [4]. In her subsequent work [5], she showed how Hecht-Nielsen’s network could be approximated using a traditional NN framework. Sprecher further contributed by deriving numerical algorithms for implementing both the internal and external functions in KART [6, 7].
Igelnik and Parikh later proposed what they termed Kolmogorov’s spline network, effectively an early form of KAN, in which both internal and external functions were represented using cubic spline basis functions [2]. A similar spline-based approach employing cubic B-splines was introduced by Coppejans [8]. Throughout the 1990s and 2000s, various other types of basis functions were explored. For instance, Polar and Poluektov employed piecewise-linear functions whose nodes were adaptively adjusted during training [9]. They demonstrated that the Kolmogorov-Arnold representation is not merely a composition of functions but can also be viewed as a specific case of a tree of discrete Urysohn operators. Moreover, they proposed a fast and computationally stable deep learning algorithm for constructing such Urysohn trees. A practical implementation of a multilayer KAN architecture, along with open-source code, was later provided by the same authors [10]. In their subsequent work [11], it was shown that a KAN model trained using the Newton-Kaczmarz method is less sensitive to the choice of the initial guess and learns significantly faster than an MLP while maintaining comparable accuracy.
@@
Despite these advancements, KANs are not without criticism. Some studies argue that KAN outperforms MLPs primarily in symbolic formula representation but falls short in tasks like computer vision, natural language processing, and audio processing [18]. Tran et al. [19] reported that despite their theoretical advantages, KANs do not consistently outperform MLPs in practical classification tasks. Additionally, their hardware implementations tend to be less efficient, with higher resource usage and latency. Sensitivity to noise can be another limitation; even minimal noise in the data can significantly degrade performance [20].
@@
The flexibility of KANs has led to explorations with various activation functions beyond B-splines, including piecewise-linear functions [9], wavelets [32], radial basis functions [33], Fourier series [34], Jacobi basis functions [35], rational functions [36], and ReLU [37]. A comprehensive comparison of activation functions used in KAN architectures is available in [38].
@@ I-B Motivation and Contributions
KAN models are traditionally trained using the backpropagation algorithm. In this work, we use the term backpropagation in its broader, commonly accepted sense – that is, as the core learning mechanism typically employed to train NNs by minimizing a loss function through gradient-based optimization. In this context, it encompasses both the computation of gradients and their use by an optimization algorithm (e.g., stochastic gradient descent) to iteratively update model parameters across layers.
However, gradient-based learning processes are prone to challenges such as local minima, flat regions, and saddle points in the loss landscape. In addition, gradient computation can be computationally demanding, particularly for deep or wide architectures, complex target functions, and large-scale datasets. These methods may also suffer from the vanishing or exploding gradient problem, which can significantly hinder convergence and stability in deep NNs.
@@
Our approach begins with utilizing fixed parameters for basis functions, determined either by data or randomly. These basis functions are then combined in multiple blocks using linear regression. The resulting block functions (activation functions) are subsequently combined through linear regression, and this iterative process is repeated across subsequent layers to form higher-level representations (such layer-wise learning – progressively improving higher-level abstractions of the input and enhancing generalization – has been employed before, albeit in a fundamentally different manner, for instance, in the greedy layer-wise unsupervised training strategies used in Deep Belief Networks [61, 62]). Combining diverse blocks corresponds to ensembling, while performing it layer by layer constitutes a multi-stacking approach. This hierarchical modeling of the target function progressively enhances accuracy at each level, eliminating the need for backpropagation.
@@
    Multi-Stacking Approach for Prediction: Our hierarchical KAN implements hierarchically multi-stacking approach to built predictions. In each layer, meta-learners combine predictions performed by weak learners (univariate models). Subsequent layers, fed by predictions from previous layers, successively refine the results, enhancing overall accuracy layer by layer.
@@
    Empirical Results for Regression Problems: We provide comprehensive empirical evidence demonstrating that our HKAN outperforms standard KAN and MLP in a range of regression problems.
@@ II Preliminary
II-A Kolmogorov-Arnold Representation Theorem
KART, also known as the superposition theorem, stands as a cornerstone in the theory of function approximation. This profound result asserts that any continuous function of several variables can be represented as a composition of continuous functions of one variable and addition.
@@ II-A Kolmogorov-Arnold Representation Theorem
KART stands as a bridge between pure mathematics and applied computational science, highlighting the potential for representing complex functions through simpler components while also illustrating the challenges in translating theoretical results into practical applications.
II-B Kolmogorov-Arnold Networks (KANs)
Paper [1] extends and modifies KART to create KANs in several key ways. While the original theorem uses a 2-layer network with a specific width in the hidden layer, KANs generalize this to allow arbitrary widths and depths, stacking multiple ”KAN layers”.
A KAN layer is defined as a matrix of activation functions $\phi_{q,p}$, where $q$ is not restricted to the theoretical limit of $2n+1$:
@@ II-B Kolmogorov-Arnold Networks (KANs)
where $x_{p}$ denotes the $p$-th input to the layer, $n_{in}$ denotes the number of inputs, and $n_{out}$ denotes the number of outputs (not restricted to $2n+1$).
@@
Furthermore, the paper provides tools for users to visualize and modify KANs, making them more interpretable and interactive. This approach takes the core idea of representing multivariate functions using univariate functions and addition, and extends it into a flexible, trainable NN architecture with theoretical guarantees and practical advantages over standard MLPs. In essence, [1] modernizes KART for use in contemporary machine learning, offering a new perspective on function approximation and NN design.
@@ III Hierarchical KAN
NN number of training samples
@@
III-A Model
HKAN is an advanced NN architecture inspired by KART. While it shares similarities with the KAN architecture, HKAN introduces unique components and a distinct training process. The architecture of HKAN is shown in Fig. 1.
@@ III-A Model
Let $\mathbf{z}^{(l-1)}$ denote the input vector to layer $l$. The first layer of HKAN receives the input vector $\mathbf{z}^{(0)}=\mathbf{x}=[x_{1},\ldots,x_{n}]^{\top}\in\mathbb{R}^{n}$. Each component of the input vector is processed by a corresponding group of blocks. Within each group, there are $n^{(l)}$ blocks, and each block applies a nonlinear projection of its input using $m^{(l)}$ basis functions (BaFs). Consider two common basis functions: the Gaussian function
@@
The BaFs within a block, denoted $g^{(l)}_{q,p,r}$$\mathbb{R}\rightarrow\mathbb{R}$, are combined linearly :
The resulting function $\phi^{(l)}_{q,p}:\mathbb{R}^{m^{(l)}}\rightarrow\mathbb{R}$ is referred to as a block function (BlF), corresponding to the inner function in (1). Fig. 2 shows the composition of a block function using BaFs.
Figure 2: Illustration of block function composition using BaFs.
Subsequently, the $n^{(l-1)}n^{(l)}$ BlFs are linearly combined by $n^{(l)}$ $h$-functions, corresponding to the outer functions defined in (1). Each $q$-th $h$-function takes as input the $q$-th BlF from every group:
The output of layer $l$ is given by $\mathbf{z}^{(l)}=\hat{\mathbf{y}}^{(l)}=[\hat{y}^{(l)}_{1},...,\hat{y}^{(l)}_{n^{(l)}}]^{\top}\in\mathbb{R}^{n^{(l)}}$, where each component $\hat{y}^{(l)}_{q}=h^{(l)}_{q}(\mathbf{z}^{(l-1)})$. This output is then fed to the next layer and processed using (5) and (6).
Following the KAN design introduced in [1], the structure of the top layer ($L$) in HKAN slightly differs from that of the preceding layers, as it contains only one BaF per input. The final output of HKAN is then produced as a linear combination of the $n^{(L)}=n^{(L-1)}$ BlFs:
$h^{(L)}(\mathbf{z}^{(L-1)})=\sum_{p=1}^{n^{(L)}}w^{(L)}_{p}\phi^{(L)}_{p}(z^{(L-1)}_{p})$ (7)
III-B Learning
HKAN learning is hierarchical and proceeds layer by layer. Below, we discuss the parameter selection/estimation process for BaFs, BlFs, and $h$-functions. The parameters of the BaFs are determined in a randomized manner (without learning), whereas for the BlFs and $h$-functions, parameter estimation is performed using linear regression.
III-B1 BaFs
[…]
@@ III-B1 BaFs
The support point method aligns the BaFs with the data distribution, thereby avoiding empty regions in the input space. In Fig. 2, the red markers represent the support points that determine the placement of the BaFs.
III-B2 BlFs
The weights of the BaFs in (5), denoted as $c^{(l)}_{q,p,r}$, are determined using the least-squares method by minimizing the sum of squared residuals:
@@ III-B2 BlFs
where $\mathbf{c}^{(l)}_{q,p}=[c^{(l)}_{q,p,1},...,c^{(l)}_{q,p,m^{(l)}}]^{\top}$, $\mathbf{y}=[y_{1},...,y_{N}]^{\top}$ and $\mathbf{G}^{(l)+}_{q,p}$ is the Moore-Penrose generalized inverse of the BaF response matrix to the $N$ training data points or their projections:
@@
III-B3 h-functions
The weights of combination (6) are calculated as:
@@
where $\mathbf{w}^{(l)}_{q}=[w^{(l)}_{q,1},...,w^{(l)}_{q,n^{(l)}}]^{\top}$, and $\mathbf{\Phi}^{(l)+}_{q}$ is the Moore-Penrose generalized inverse of the BlF response matrix to the projections of $N$ training data points:
@@
Table II summarizes the HKAN parameters and the stages of the hierarchical learning process.
TABLE II: Overview of the HKAN parameters and the sequential stages of its hierarchical learning process.
Layer 1
HKAN component BaFs BlFs $h$-functions
Parameters $\mu^{(1)}_{q,p,r}$ $c^{(1)}_{q,p,r}$ $w^{(1)}_{q,p}$
Number of parameters $\cdot n^{(1)}\cdot m^{(1)} nn$$\cdot n^{(1)}\cdot m^{(1)}$ $n\cdot n^{(1)}$
Learning method None; fixed parameters Least squares with target $y$ Least squares with target $y$
(random/data-driven)
Number of least-squares runs - $n\cdot n^{(1)}$ $n^{(1)}$
Output $g^{(1)}_{q,p,r}$ – nonlinear randomized $\phi^{(1)}_{q,p}$ – linear combinations $h^{(1)}_{q}$ – linear combinations
projections of inputs $x_{p}$ of BaFs in each block of selected BlFs
Layer 2
HKAN component BaFs BlFs $h$-functions
Parameters $\mu^{(2)}_{q,p,r}$ $c^{(2)}_{q,p,r}$ $w^{(2)}_{q,p}$
Number of parameters $n^{(1)}\cdot n^{(2)}\cdot m^{(2)}$ $n^{(1)}\cdot n^{(2)}\cdot m^{(2)}$ $n^{(1)}\cdot n^{(2)}$
Learning method None; fixed parameters Least squares with target $y$ Least squares with target $y$
(random/data-driven)
Number of least-squares runs - $n^{(1)}\cdot n^{(2)}$ $n^{(2)}$
Output $g^{(2)}_{q,p,r}$ – nonlinear randomized $\phi^{(2)}_{q,p}$ – linear combinations $h^{(2)}_{q}$ – linear combinations
projections of inputs $z^{(1)}_{p}$ of BaFs in each block of selected BlFs
Layer 3 (top)
HKAN component BaFs BlFs $h$-functions
Parameters $\mu^{(3)}_{q,p,r}$ $c^{(3)}_{q,p,r}$ $w^{(3)}_{p}$
Number of parameters $n^{(3)}\cdot m^{(3)}$ $n^{(3)}\cdot m^{(3)}$ $n^{(3)}$
Learning method None; fixed parameters Least squares with target $y$ Least squares with target $y$
(random/data-driven/even distrib.)
Number of least-squares runs - $n^{(2)}\cdot n^{(3)}$ 11
Output $g^{(3)}_{p,r}$ – nonlinear randomized $\phi^{(3)}_{p}$ – linear combinations $h^{(3)}$ – linear combinations
projections of inputs $z^{(2)}_{p}$ of BaFs in each block of BlFs